### v6

# Enhancing Efficiency in Multidevice Federated Learning through Data Selection

Abstract.
UbiquitousTo be presented in the 10th ACM/IEEE Symposium on Edge Computing (SEC’25), December 2025, Arlington, VA, USA. wearable and mobile devices provide access to a diverse set of data. However, the mobility demand for our devices naturally imposes constraints on their computational and communication capabilities. A solution is to locally learn knowledge from data captured by ubiquitous devices, rather than to store and transmit the data in its original form. In this paper, we develop a federated learning framework, called Centaur, to incorporate on-device data selection at the edge, which allows partition-based training of a deep neural nets through collaboration between constrained and resourceful devices within the multidevice ecosystem of the same user. We benchmark on five neural net architecture and six datasets that include image data and wearable sensor time series. On average, Centaur achieves $\sim$19% higher classification accuracy and $\sim$58% lower federated training latency, compared to the baseline. We also evaluate Centaur when dealing with imbalanced non-iid data, client participation heterogeneity, and different mobility patterns. To encourage further research in this area, we release our code at github.com/nokia-bell-labs/data-centric-federated-learning.
Key words and phrases:
Federated Learning, Constrained Devices, Data Selection, Partition-Based Training, Edge Intelligence, On-device AI
1. Introduction
Figure 1. The latency of federated training versus classification accuracy on test dataset. We compare Centaur with standard federated training that only trains the classifier on ubiquitous constrained devices (UCDs) without data selection and partition-based training. (Top plot) Four models were tested on image data from CIFAR-10, and (bottom plot) a model was tested on three datasets of wearable sensor time-series.
With the growing trend of utilizing ubiquitous devices in personal and industrial environments, federated learning (FL) enables the discovery of new patterns in personal data by training deep neural networks (DNNs) on data captured by various users in a more private manner floriginal; flsurvey. Especially, in a multidevice ecosystem, the owner can seamlessly share data across the devices within their trusted environment while protecting data privacy cho2022flame. This allows us to utilize the capabilities of multiple devices to participate in the FL process. However, wearable and ubiquitous devices have limited computing resources, little on-device storage, and inconsistent network connectivity.
Most of the existing FL frameworks do not account for constraints of personal devices and instead assume either modern smartphones hard2018federated; niu2020billion, or edge devices with DNN accelerators cho2022flame; zhang2022federated, as target client devices. Although smartphones and accelerators are less powerful than cloud machines, they are nevertheless equipped with mobile GPUs, possess gigabytes of runtime memory, and have fairly stable connectivity to a central server, which simplifies the requirements of DNN training yang2021flop; lin2020ensemble.
In this paper, we propose a novel FL framework to include ubiquitous constrained devices (UCDs), i.e., wearable devices such as earbuds, glasses, or rings, or distributed edge devices such as environmental cameras or industrial sensors. We believe UCDs are becoming the primary data-producing devices for both individuals and industries, due to various on-device sensors, such as cameras or inertial measurement units. These devices offer a wealth of spatiotemporal data that is frequently absent when focusing solely on smartphones or plugged-in devices. For example, people often opt for wristbands, smart rings, or earbuds over carrying a smartphone in their hand or pocket while engaging in outdoor, industrial, or sports activities seneviratne2017survey; mo2021adapting, and sensors distributed across farms or forests remotely monitor crops and livestock to detect hazards papst2021exploring. The main motivation of our work is the fact that the abundance and diversity in data captured by ubiquitous devices can facilitate learning more effective models for emerging applications in personal and industrial environments.
There are several significant challenges. Usually, distributed or wearable UCDs use an access point (AP), like a smartphone or a router, as a relay to connect to a central server in the cloud. Data captured by a UCD is communicated to the AP using communication protocols like Bluetooth Low Energy (BLE). In outdoor settings under mobility, UCDs may not always be connected to the companion AP, thus inaccessible to the central server. Beside this, the constrained data storage and computing resources will restrict the size and functionality of DNNs that can be trained locally on UCDs. A naive solution is to transfer all data from UCDs to their respective APs and perform training of DNNs on APs. Yet, storing all data could exceed the device’s memory, especially when UCDs disconnect from their AP. To demonstrate these points, in §3, we conduct a motivating study on a real testbed to identify the essential design requirements that can enable the deployment of FL on resource-constrained devices. To tackle these challenges effectively, we believe FL solutions relying on on-device training should dynamically distribute data and computations between UCDs and their companion APs.
Table 1. Comparison with related work. (1) Model Partitioning splits the model into an encoder and a classifier. (2) Data Selection trains on a subset of available data. (3) On-device Training considers backpropagation on constrained devices. (4) Spatiotemporal Coverage includes mobility patterns of portable devices.
Model Partitioning Data Selection On-device Training Spatiotemporal Coverage
Multitier FL abad2020hierarchical; liu2020client; abdellatif2022communication; cho2022flame; gupta2018distributed × × × ×
Split FL diao2020heterofl; zhang2021fedzkt; zhang2021parameterized; li2021hermes; dai2022dispfl; bibikar2022federated × × ×
On-device Training lee2019intermittent; islam2019zygarde; fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; lin2022device; kopparapu2022tinyfedtl × ×
Centaur (ours)
In §4 and Figure 3, we introduce Centaur: a federated learning framework that orchestrates local training among UCDs and APs by integrating on-device data selection with partition-based model training.
(A) Considering computation constraints, we initialize the DNN in two partitions: an encoder (i.e., feature extractor) to be only trained on APs, followed by a lightweight classifier to be trained on both APs and UCDs (§4.1). This allows Centaur to utilize all the data available on UCDs for fine-tuning the DNN while occasionally storing and transmitting a portion of the data to the APs to re-train the entire DNN and adjust to the runtime variability in data distributions gupta2018distributed; thapa2022splitfed; weiss2016survey.
(B) Considering memory and connectivity constraints, we perform data selection jiang2019accelerating by analyzing the training loss and the gradients norm of the classifier part, to decide which data points captured by UCDs contribute more to the training of which part of the DNN. Through data selection, data points are categorized as either of (i) discarded if they have very low loss values, (ii) kept locally on the UCD to train the classifier part if their loss values or gradients norm are not very high, and (iii) transmitted to the AP to train both encoder and classifier part if they cause high values for both loss and gradients norm (§4.2). We assume that each pair of UCD and AP belongs to the same client, thus transmitting data from a UCD to its AP does not violate the clients’ privacy.
(C) We benchmark on five DNNs architectures and six datasets of two different modalities, and compared to existing FL alternatives. Our results show that Centaur saves more bandwidth by reducing the communication cost associated with offloading samples from a UCD to the AP, and also reduces the latency of training (and accordingly the energy consumption) on both UCD and AP. As a prime example of our experimental results, Figure 1 shows the training latency versus classification accuracy of performing FL for two different tasks: (1) four benchmark DNNs, with their encoder pre-trained on ImageNet deng2009imagenet, and then trained for CIFAR10 krizhevsky2009learning image classification with 100 FL clients, and (2) a benchmark ConvNet chang2020systematic, trained from scratch, on three human-activity recognition datasets with 20 FL clients. In both figures, we compare Centaur with the baseline of standard FL on UCDs without our implemented strategies. Centaur, through data selection and partition-based model training, achieves up to $19\%$ higher accuracy and $58\%$ lower latency, on average.
The key contribution of our work is a better integration of ubiquitous constrained devices into federated learning by leveraging the advantages of multi-device ecosystems. We achieve this by introducing a customizable data selection scheme and a partition-based training approach, which collectively reduce computational and communication costs while enhancing model accuracy. Our experiments demonstrate that Centaur effectively minimizes storage requirements and training time through efficient data selection. Furthermore, our analysis accounts for the real-world spatiotemporal mobility of edge devices in FL. Empirical evaluations, covering cost, data imbalance, participation heterogeneity, and connection probability, show that Centaur consistently delivers higher efficiency, achieving improved accuracy at lower costs across diverse scenarios.
2. Related Work
Table 1 compares Centaur with major prior work on different aspects of multidevice FL. Compared to the standard edge–server FL, edge–access point–server FL liu2020client; abad2020hierarchical considers a middle layer, e.g., a cellular base station, to orchestrate FL training across clients and a central server. Clients can participate with several devices cho2022flame; diao2020heterofl; zhang2021fedzkt, which achieve better trade-offs between accuracy and bandwidth consumption. However, these works assume the local training of the entire model on edge devices, without any resource limitations. They also do not utilize the potential of resource-rich devices to assist a resource-constrained device during training. They use APs only for aggregation and communication, which makes FL impractical in scenarios involving UCDs.
Other works consider inference only (i.e., forward pass) fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas, sparse training li2021hermes; bibikar2022federated, or heterogenous architectures across clients of different capabilities kang2023nefl; park2024federated. Sparse training or pruning requires computational resources (especially memory) that UCDs do not have. Few works propose solutions for full on-device training (i.e., forward and backward pass) lin2022device, but not for FL scenarios as they often make impractical assumptions like the availability of large datasets for network architecture search liberis2021munas, or unconstrained memory for training the full model diao2020heterofl; zhang2021fedzkt. Transfer learning can be used on UCDs to fix the encoder architecture, which in turn restricts the extraction of newer features, and only trains the last layer for the personalized training zhang2021fedzkt.
Without utilizing the power of APs, training the entire model on UCDs requires a large sample size to learn generalized features li2021hermes; bibikar2022federated. Simply considering different sub-models diao2020heterofl; kang2023nefl; park2024federated requires more sophisticated aggregation strategies and shallower models for resource-constrained devices, which may not be capable of learning the complex patterns from the data. Recent works enable training on micro-controllers with limited memory, such as TensorFlow Lite Micro david2021tensorflow or Tiny Training Engine lin2022device. These works do not consider FL as a potential use case and do not account for the advantages of a multidevice ecosystem.
3. A Motivating Study
As depicted in Figure 3, a ubiquitous constrained device (UCD) is a small, low-power portable device with restricted hardware and connectivity resources. Examples include smart rings, earbuds or glasses (personal wearables) and tiny environmental or industrial sensors. Each UCD connects to the user’s personal access point (AP), usually over a low-power link such as Bluetooth Low Energy. An AP is a resourceful device that has much better hardware and connectivity capabilities than its corresponding UCD (for instance, a smartphone, home router, or gateway). We assume the AP is always part of the user’s device ecosystem that allows the UCD to transmit data to the AP without violating the user’s privacy. We generally assume one AP serves multiple UCDs (e.g., a person’s phone might pair with a smartwatch, earbuds, and other sensors simultaneously). Indeed, modern homes routinely host on the order of 20–30 connected devices, so a single AP often handles many UCDs.
A server is a central entity owned by a service provider that orchestrates FL over several clients. Considering a deep neural network architecture, an encoder ($E$) is the first part of the DNN that extracts the features from input data. A classifier ($C$) is the second part of the DNN, which performs the final classification using the encoded features. We consider supervised learning tasks where $\mathbf{X}$ denote the input data and $\mathbf{y}$ denote its label.
Setup. To identify the key design choices, we create a testbed with four RaspberryPi 4 Model B as edge devices velasco2018performance; drakopoulos2019real; all running the 64-bit Raspbian OS with total secondary storage of 3232GB. We set up four clients with varying primary memories: C1, C2, C3, and C4 having 11GB, 22GB, 44GB, and 22GB, respectively. Indeed, Raspberry Pi devices are more capable than many wearables; however, they serve as a widely adopted, reproducible proxy for constrained clients cho2022flame, enabling controlled CPU/memory benchmarking while we contextualize our findings with real-world wearable hardware specifications and multitasking considerations. For analyzing the CPU and memory consumption of the clients, we use logs of /proc/stat and free operating-system calls, respectively. We use MobileNetV3 howard2019searching, pre-trained on ImageNet, as the encoder followed by two fully-connected layers with an intermediate drop-out layer. Using CIFAR10, we measure the resources and time needed for both on-device training (per epoch) and inference. We implement FL with Flower beutel2022flower and use a desktop computer as the server, as shown in Figure 2a.
Figure 2. On four RaspberryPis, we measure memory consumption of running FL for when (i) only the classifier is trained, (ii) the entire model is trained, and (iii) the entire model runs inference only.
[…]
Key Observations. We run FedAvg floriginal for $100$ rounds with four RaspberryPi devices, considering two scenarios: (1) when only the classifier part is trained, and (2) when the entire model (encoder and classifier) is trained. When running our experiments for the first scenario, we observed that the standard FL runs successfully over all four clients. However, for the second scenario, client C1 with 1GB of primary memory always failed to participate in the FL process. This is primarily due to the significant rise in the CPU and memory consumption when the entire model (both encoder and classifier) are trained on the device (see Figure 2).
Notably, 1GB of primary memory is still significant in terms of memory capability, and many UCDs may have significantly less memory than this. Therefore, it is clear that in such cases, it is much more challenging to deploy standard FL over UCDs. Moreover, training only the classifier consumes far less resources and is even comparable to the resource usage required for a forward pass. We noticed a similar pattern for time consumption, where training the classifier only is $50\times$ time efficient compared to training the entire model. One alternative solution to decrease the overall training time is to reduce the training data size by randomly discarding samples. However, we observe that such an approach comes at a cost of significantly reduced accuracy, which is detrimental to the overall system.
4. Method
In Figure 3, we introduce Centaur, our federated learning framework. Here, we elaborate on each component in detail.
[…]
Since full backpropagation involves a forward pass, the storage of all intermediate activations, and a backward pass through every layer to compute gradients, it is the most compute- and memory-intensive phase of model training. Centaur mitigates this by freezing the deep encoder and training only the lightweight classifier head on UCDs. Each device performs a single forward pass through the fixed encoder and backpropagates only through the shallow fully-connected classifier, which has orders of magnitude fewer parameters. This design significantly reduces the on-device workload. Our motivationa study in §3 confirmed this trend, with classifier-only training running several times faster and using far less memory than updating the entire network. Thus, Centaur enables effective training on local data while keeping UCD training lightweight and efficient.
4.1. Model Initialization
To partition a DNN into an encoder $E$ and a classifier $C$, one option is to use pre-trained benchmark DNNs; considering numerous well-trained models. Usually, DNN architectures from a candidate space of pre-trained models have already been well developed, with efficient quantization and compression capabilities, to be deployed on UCDs for inference fedorov2019sparse; banbury2021micronets; xu2022etinynet; liberis2021munas; li2021hermes. We aim to utilize the full capability of UCDs for on-device training of the classifier $C$. Thus, we need to design the classifier such that it can run on the limited memory and computation power available on these devices. To achieve these, we can iteratively look into candidate architectures that fit in the limited memory available on these UCDs. A robust mechanism in this direction is to obtain a history of memory usage of a particular device, which reflects the typical memory availability on the UCD.
To make the selection process explicit, we consider a set of lightweight architectures $\{m_{k}\}$: a small set of widely used mobile convolutional models (e.g., SqueezeNet and MobileNet-style variants), alongside fully connected networks with varying depth (1–3 layers) and width (16–128 neurons). These architectures are chosen because they span the typical memory–compute trade-off space encountered in edge devices while still supporting on-device gradient updates. Among the feasible candidates that satisfy the device’s resource constraints, we select the smallest model that achieves an accuracy within a small tolerance of the best-performing candidate. This strategy effectively balances efficiency and performance, ensuring that the chosen classifier maximizes on-device feasibility without incurring a significant loss in learning effectiveness.
Formally speaking, the iterative process can be viewed as an optimization problem described as
$C=\operatorname*{argmin}_{\{m_{1},\dots,m_{K}\}}\;\mathbb{P}(m_{k})\,\mathbb{B}\penalty 10000\ \penalty 10000\ \text{s.t.}\penalty 10000\ \penalty 10000\ \mathbb{P}(m_{k})\,\mathbb{B}\leq\mathbb{M}\penalty 10000\ \text{and}\penalty 10000\ \text{acc}(m_{k})\geq\tau^{*},$ (1)
[…]
21Inputs: (1) $\mathcal{A}$: the set of AA clients each having two devices, one UCD and one AP, (2) $D^{k}$: local dataset for each client $a\in\{1,\dots,A\}$, (2) $f_{E}$: the encoder part with parameters $W_{E}$, (3) $f_{C}$: the classifier part with parameters $W_{C}$, (4) $R$: the total number of FL rounds.
30.1cm
54 for $r\in\{1,\dots,\nicefrac{{R}}{{2}}\}$ rounds do
7 6 $K$ clients $\Leftarrow$ randomly select $K$ clients from a total of AA, using uniform sampling
8 0.1cm
9  — Training on UCDs —
11 10 for $k\in\{1,...,K\}$ do
13    12 $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D^{k}$ (all the samples in the local dataset of client $k$ UCD)
15    14 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
17    16 $D_{N}^{k},D_{C}^{k},D_{M}^{k}\Leftarrow$ based on $\ell^{k}$, perform data selection as detailed in Section 4.2
19    18 for all $(\mathbf{X}^{k}_{j},\mathbf{y}^{k}_{j})\in D_{C}^{k}$ do
21       20 $g^{k}_{C}={\partial\ell^{k}}/{\partial W_{C}}$ (compute $f_{C}$’s gradients on $D_{C}^{k}$ to update $f_{C}$)
23       22 $W_{C}^{k}=Optimizer(W_{C}^{k},g^{k}_{C})$ (update $f_{C}$’s parameters using gradients)
25       24 $D^{+}\Leftarrow$ based on $g^{k}_{C}$, perform data selection as detailed in Section 4.2
27       26 $D_{M}^{k}=D_{M}^{k}\cup D^{+}$ (enhance $D_{M}^{k}$ by high-value gradients data points)
28  — Server Side Aggregation on the Classifier Part —
30 29 $W_{C}\leftarrow\frac{1}{K}\sum_{k\in K}W_{C}^{k}$
32 31 $W_{C}^{k}\leftarrow W_{C}$ (update all clients with new $W_{C}$)
33 0.1cm
34  — Training on APs —
36 35 for $k\in\{1,\dots,K\}$ do
37
39    383pt $\mathbf{X}^{k},\mathbf{y}^{k}\leftarrow D_{M}^{k}$ (all selected samples for training $f_{E}$ on client $k$ AP)
41    40 $\ell^{k}=\mathcal{L}(f_{E}\circ f_{C}(\mathbf{X}^{k}),\mathbf{y}^{k})$ (forward pass to compute per-sample loss values)
43    42 $g^{k}={\partial\ell^{k}}/{\partial(W_{E},W_{C})}$ (compute gradients on $D_{M}^{k}$ to update $f_{E}$ and $f_{C}$)
45    44 $W_{E}^{k},W_{C}^{k}=Optimizer(W_{E}^{k},W_{C}^{k},g^{k})$ (update all parameters using gradients)
46  — Server Side Aggregation on both Encoder and Classifier Parts —
48 47 $\{W_{E},W_{C}\}\leftarrow\frac{1}{K}\sum_{k\in K}\{W_{E}^{k},W_{C}^{k}\}$
50 49 $\{W_{E}^{k},W_{C}^{k}\}\leftarrow\{W_{E},W_{C}\}$ (update all clients with the new aggregated model)
51 return $\{W_{E},W_{C}\}$
Algorithm 1 Centaur: Multidevice Federated Learning via Partition-based Training and Data Selection
4.2. Data Selection
We determine the importance of each data sample, along the forward pass, to decide whether to further perform (the more costly) backward pass on this sample or not. Such data selection satisfies design requirements by improving both storage efficiency and training efficiency. We use a combination of (i) the loss value and (ii) the norm of last-layer gradients to measure the sample’s importance.
4.2.1. Loss-based Selection.
[…]
$\begin{cases}\mathcal{P}_{N}(\ell_{r+1})=1-[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\alpha}\\ \mathcal{P}_{M}(\ell_{r+1})=[\text{CDF}^{\ell}_{r}(\ell_{r+1})]^{\beta}\end{cases}$ (2)
[…]
4.2.2. Gradient-based Selection.
For samples that are not selected through $\mathcal{P}_{N}$ and $\mathcal{P}_{M}$, and to avoid consuming extra resources of UCDs, we can further compute the gradients of the classifier’s last layer. The norm of these gradients gives us a useful hint about the sample’s importance while requiring much less computation than computing all layers’ gradients katharopoulos2018not. The last layer’s norm at round $r$ is produced during training of the classifier by $g_{r}=\frac{\partial\ell_{r}}{\partial W_{r}}$, where $W_{r}$ is the weights of the classifier $C$’s last layer at round $r$. Following the same idea, we derive a CDFg for gradient norm values and build a queue to keep track of the computed norm of gradient values. Samples with larger gradients have larger impacts on the model’s weights; thus, we also keep these samples for training the entire model $M$ as $E$ might learn new “features” from them. Thus, the probability of adding a sample to $D_{M}$ at round $r+1$ is defined as:
$\mathcal{P}_{M}^{+}(\|g_{r+1}\|)=[\text{CDF}^{g}_{r}(\|g_{r+1}\|)]^{\gamma}$ (3)
where $\gamma$ is used for customizing the selection rate. Samples with high norm are selected with $\mathcal{P}_{M}^{+}(\|g_{r+1}\|)$ and then added to data $D_{M}$. Our dynamic strategy, in combining loss and last-layer gradient norm, enables Centaur to achieve a better trade-off between computation cost and selection performance on UCDs. Moreover, one may substitute this module with a different data selection technique that could also be dependent upon the specific use case.
While our proposed data selection strategy is primarily empirical, it is motivated by the theoretical principles of importance sampling in deep learning. Prior work has shown that sampling data points in proportion to their loss values or gradient magnitudes can approximate the optimal sampling distribution that minimizes the variance of stochastic gradient estimates and thus accelerates convergence katharopoulos2018not; jiang2019accelerating. In particular, high-loss samples tend to carry larger gradient norms and contribute more substantially to the update direction, while samples with small losses or small gradient norms contribute less. Our combination of loss value and last-layer gradient norm therefore aligns with these theoretical insights: it adaptively emphasizes samples expected to produce high-informative gradients, balancing computational cost with learning efficacy. Nonetheless, we acknowledge that our current formulation does not provide a formal convergence guarantee. Developing theoretical guarantees through variance-reduction or convergence-rate analysis is a promising direction for future work.
4.3. Partition-based Training and Aggregation
The results of §3 highlight why we freeze the encoder and only train the classifier on UCDs. Training just the classifier consumed far less memory and CPU, comparable to running an inference pass, and ran much faster than full-model training. Since most UCDs have far less memory than even a RaspberryPi (e.g. a few MB of storage or only a few hundred KB of SRAM), full-model training on UCDs is generally infeasible. By freezing the encoder on UCDs, Centaur aims to reduce the on-device computation and memory requirements, allowing each UCD to train only the compact classifier locally. The more powerful AP (smartphone or router) can then periodically update the encoder on the offloaded samples, so that the overall model can still learn richer features even though each UCD only fine-tunes the classifier part.
Therefore, our training is conducted on both UCDs and APs. While UCDs only train the classifier $C$ on data $D_{C}$, APs train the complete model $M$ on data $D_{M}$. Such a procedure allows the model to be (partially) updated when UCDs are offline, in addition to the full updates when UCDs have a connection to the internet. This better utilizes UCDs’ spatiotemporal richness and consequently improves the training efficiency in our design requirements.
Algorithm 1 shows partition-based training and aggregation in Centaur, running $\nicefrac{{R}}{{2}}$ rounds, considering that UCD training and AP training proceed iteratively in one round. All training is conducted on the client’s devices either on UCDs or APs. Then with all updated weights, the server performs Federated Averaging (FedAvg) floriginal to obtain one aggregated global model. The new global model needs to be distributed to all clients UCDs and APs such that (i) the training on AP has an updated classifier, and (ii) the training on UCD in the next round has both an updated encoder and classifier.
Centaur is the first FL framework to simultaneously combine model partitioning, data selection, and on-device training. Our approach of partitioning the DNN into encoder and classifier, selecting a subset of data on-device, and performing lightweight on-device updates, has not been addressed jointly in earlier work. In contrast, existing hierarchical FL schemes either train full models on each device or only aggregate updates at the AP without splitting the model.
5. Experimental Setup
Datasets. We use six commonly used datasets. CIFAR10 and CIFAR100 krizhevsky2009learning that both include 50K training samples and 10K test samples, with 10 and 100 classes, respectively. EMNIST cohen2017emnist that includes 112K training samples with 47 classes. UCIHAR Anguita2013APD is a widely used dataset of 30 users performing 6 daily activities; data from the accelerometer and gyroscope sensors were collected by a smartphone worn on the waist. Data from 21 users is used for training, and that of the other 9 users for testing purposes. MotionSense malekzadeh2019mobile also includes accelerometer and gyroscope data from 24 users with a smartphone in the pocket of the trousers who performed 6 activities in 15 trials. We use as test data one trial session for each user and as training data the remaining trial sessions (e.g., one trial of “walking” of each user is used as test data and the other two trials are used as training). PAMAP2 reiss2012introducing dataset contains data of 13 different physical activities, performed by 9 subjects wearing 3 devices and a heart rate monitor. We use the training-test split provided by the PersonalizedFL library PersonalizedFL.
Models. For image classification datasets, we select four DNNs commonly used in mobile/edge-oriented literature: (1) EfficientNet-v2 tan2021efficientnetv2, (2) MobileNet-v3 howard2019searching, (3) MNASNet tan2019mnasnet, and (4) ShuffleNet ma2018shufflenet, with 5.3M, 2.2M, 2.5M, and 1.4M number of parameters respectively. For human-activity recognition datasets, we borrow the ConvNet architecture proposed in chang2020systematic. All these models consist of an encoder made of multi-layer CNNs followed by a classifier.
FL Settings. We use Flower beutel2020flower, a customizable open-source FL framework (Python v3.7, Ray v1.11, and Torch v1.12). We run all simulations on a server with 80 Intel Xeon(R) E5-2698 CPUs, 8 Tesla V100 GPUs (16GB), and 504GB system RAM. For all simulations, we use 100 communication rounds, where at each round, UCDs and APs successively train the model. Unless specified otherwise, we consider 100 clients (each having one AP and one UCD). We randomly sample 10% of the clients for training in each round in our initial experiments and later show that our results hold when the client participation ratio is gradually increased to 100%. Data ($D$) are partitioned using Latent Dirichlet Allocation (LDA) blei2003latent; beutel2022flower without resampling (LDA-Alpha=1000). We refer to prior research blei2003latent about details of LDA’s generative process. After such partition, each client owns $\nicefrac{{|D|}}{{100}}$ local samples.
For each client, we randomly select half of the samples and use them for local training of the model. For both APs and UCDs, we consider by default the number of epochs to be 3 with a batch size of 64. The other half of the samples are available with more offline time to simulate the further collection of extra data. APs and UCDs also train more epochs on the extra data while offline. In this setup, we define every unit offline time contributing to one more epoch of training on $20\%$ of these extra local samples.
To set the hyperparameters for Eqs (2)-(3), we run an experiment on CIFAR-10 with MobileNet-V3. We evaluate the trade-off between test accuracy and the workload imposed by data selection on both UCDs and APs. This experiment can be executed by the server as a preparatory step before deploying Centaur, using small representative datasets that approximate the clients’ data distribution. The results in Figure 4 show that performing data selection reduces latency and energy consumption on UCDs by over 80%. Although this approach increases the workload on APs, this is acceptable given their higher computational capacity, since we observe a substantial improvement in model accuracy. In our experiments, we set hyperparameter $\alpha,\beta,\gamma$ as $5,3,0$, respectively, as this configuration represents the fair balance between accuracy and resource efficiency.
[…]
Baselines. (1) AP Training. We consider UCDs as data collection apparatuses only and therefore, UCDs upload collected and stored data samples to their connected APs, so that the training only happens on APs. In this case, the complete model (both the encoder and classifier) is updated because APs in general are considered to have sufficient resources. (2) UCD Training. UCDs do not upload the collected data samples. Instead, they conduct training on the data locally. However, as UCDs are typically resource-constrained, only the classifier is considered to be trained on-device. This means that the encoder part is considered to be frozen in this setting.
Mobility Model. In our case, the idea of the mobility model signifies how the mobility patterns of the UCDs (or users) impact the connectivity of the UCD to its AP and eventually to the internet. We define an exclusive Online Association Matrix $\Omega$, which is a binary matrix representing the user’s exclusive location at a given time instance. Mathematically, $\Omega$ is a binary matrix $\mathcal{T}\times\mathcal{S}$ where $\mathcal{T}$ and $\mathcal{S}$ represent the temporal and spatial granularity of $\Omega$ marix, respectively. For example, the rows may represent the time zones of the day (early morning, morning, afternoon, evening, and late night), during which the user moves between three locations like home, office, and a public park, represented by the columns of $\Omega$. Furthermore, we ensure that each row’s sum equals unity, which means the user is present exclusively at a unique location in a given temporal instance. Finally, to simulate the connectivity patterns, we define a connectivity matrix $\lambda$, a single row-vector of dimension $\mathcal{S}$, denoting the connectivity probability across different locations. More specifically, for evaluating Centaur during mobility, we generate the global connectivity matrix $\lambda$ and the generate matrix $\Omega$, both chosen uniformly at random, for each user to simulate the mobility scenario.
Figure 5. Accuracy of the best classifier for four different encoders, trained on CIFAR10, CIFAR100, and EMNIST. Dash lines ( ) above depict the upper-bound accuracy when UCD devices have no resource and connectivity constraints.
6. Evaluation Results
We present the results when running Centaur compared to conventional FL training methods and compare the approaches using different metrics. We also analyze the impact of data/participation heterogeneity and spatiotemporal coverage on training methods.
6.1. Metrics
In addition to the accuracy of the trained global model on a held-out test set we evaluate the performance of our FL framework, compared to other FL alternatives, using the following metrics.
1) Accuracy. This is the classification accuracy of the trained model on a test set hosted by the central server. For a model obtained at the end of each training round, the test accuracy is computed as
$acc\equiv\frac{1}{T}\sum_{i=1}^{T}\mathbb{1}\big(\operatorname*{argmax}\big(f_{E}\circ f_{C}(\mathbf{X}_{i})\big)=\mathbf{y}_{i}\big),$
where test set has $T$ pairs of $(\mathbf{X}_{i},\mathbf{y}_{i})$ and $\mathbb{1}(\mathrm{C})$ denote the indicator function that outputs 1 if condition $\mathrm{C}$ holds.
[…]
Figure 6. Centaur performance vs. other baselines for different sizes of classifiers: small, medium, and large. (Top) image classification, and (bottom) HAR datasets.
3) Bandwidth. As the model size (i.e., the encoder or the classifier) can be different in each FL round, we use fvcore to count the number of parameters that are communicated in each round. Based on the model size and number of communications among UCDs, APs, and the server, we compute the amount of bandwidth that is consumed. Also, we count the number of sample points that are uploaded from UCDs to APs.
Figure 7. Computation workload of model training on usage-constrained devices (Top Left) and access points (Top Right), and Communication workload of model/data sharing of usage-constrained devices (Bottom Left) and access points (Bottom Right) when achieving specific test accuracy, reported on MobileNetv3 and CIFAR10.
4) Latency. The latency usually has a linear relation with MAC operations, due to the lack of specialized accelerators liberis2021munas; lin2022device. Thus, we estimate the latency of model training based on the processor’s frequency as - $\text{Latency}=c*\frac{\text{Total MACs}}{\text{Processor Frequency}}$. The ratio between MAC operations and the processor’s instructions, $c$, is typically between $1$ and $2$ based on specific instruction sets/compilers. For simplicity, we assume that each MAC operation translates to two instructions in an MCU (i.e., $c=2$). Notice that existing processors (e.g., Intel’s Load Effective Address) complete one MAC in one instruction x86assembly), and this only scales the experimental results and does not change the conclusions. Finally, communication latency is estimated based on the total amount of data needed to be transmitted, divided by the up-link speed or the down-link speed of devices.
5) Energy. This is the total execution time multiplied by the processor’s consumed power per unit time. The energy consumption of communication can also be calculated based on the total time of transmitting data multiplied by the transmitters’ power per unit time. We set the values for UCDs and APs in our simulations as follows. We assume a typical AP to have a CPU frequency of 2GHz, storage capacity of 4GB, power capacity of 1.5mW/MHz with an uplink speed of 10Mbit/s, downlink speed 100Mbit/s, and communication energy of 10W. For a typical UCDs, we assume a CPU frequency of 100MHz, storage capacity of 5MB, power capacity of 0.05mW/MHz with an uplink speed of 2Mbit/s, downlink speed of 2Mbit/s, and communication energy of 0.0001W.
6.2. Model Accuracy
We use the CNN backbone of the four benchmark models described in §5 as the encoders. We examine three classifiers: small that is only one fully connected (FC) layer of size $z$ (number of classes), medium that has two FC layers of size 64 and $z$, and large that has two FC layers of size 128 and $z$.
We permute the four encoders and three classifiers and report the test accuracy of Centaur compared to the two other baselines: AP training and UCD training. In Figure 5 we report the highest test accuracy for each encoder across all three classifiers for three datasets. Results show that Centaur outperforms AP training by $0.53\%\sim 40.15\%$ and UCD training by $0.45\%\sim 17.78\%$, depending on the settings. There are several settings in which all FL training methods cannot reach a good accuracy, e.g., NASNet and ShuffleNet on CIFAR100, probably because of their relatively small model sizes compared to data complexity. Centaur still achieves better performance compared to the baselines.
In Figure 6 we present the accuracy of different classifiers for each encoder. In the top plot, the results of training on CIFAR10 and CIFAR100 with an accuracy higher than $30\%$ are presented for better visibility. In the bottom plot, we report the results of UCIHAR, MotionSense, and PAMAP2. The results show that Centaur outperforms both UCD training in all classifiers’ sizes. The test accuracy also tends to be similar across small, medium, and large classifiers. In addition, the classifier’s sizes may have less impact on more sophisticated encoders (e.g., EfficientNet and MobileNet), which is also observed in a previous work mo2021ppfl. This may be because an appropriate encoder already produces high-quality features for unseen data that are easy to learn (e.g., CIFAR10), and in such a case, classifier sizes do not make any difference in test accuracy. Similar patterns are observed in HAR datasets. Notice that for HAR datasets, there is currently a lack of publicly accessible pre-trained encoders to start FL with. As a result, we can observe a more significant performance gap compared to image classification, particularly for more complex datasets, e.g., PAMAP2. This lack of pre-trained encoders can also be the reason that, in HAR datasets, the performance of Centaur and AP are almost the same.
These results emphasize that with a properly pre-trained backbone, continual re-training of the encoder on APs is neither necessary nor desirable. In transfer learning, it is standard to fix, or only occasionally fine-tune, the encoder and update the lightweight classifier head to adapt to local data. In Centaur, we follow this principle: UCDs perform frequent classifier-only updates to rapidly personalize to local data, while APs periodically update the encoder to correct drift in feature extraction. This approach achieves competitive accuracy with lower communication and computation costs: frequent classifier updates capture personalization and mobility, while periodic encoder updates maintain convergence and generalization. Future work can explore how varying update frequencies affect convergence trade-offs.
Figure 8. Cost linkage analysis among MAC operations, latency, and energy consumption.
6.3. Efficiency
Cost-Accuracy Trade-off. Here, with CIFAR10, MobileNet-v3 as the encoder, and a medium size for the classifier, we compute the test accuracy for a range of MAC, Bandwidth, Latency, and Energy budgets. The average size of each sample of CIFAR10 is $30KB$. Figure 7 (top) shows on both UCD and AP and across accuracy levels, Centaur achieves lower MAC and Latency than both UCD training and AP training. Figure 7 (bottom) shows that communication Bandwidth and Latency of Centaur is almost the same as AP training; however, Centaur can achieve a higher test accuracy along with more training steps when consuming more resources. Centaur uses much more communication resources than UCD training, as Centaur transmits both the classifier, although it is much lighter than the encoder, and data samples. However, the accuracy of UCD training cannot go further than $83.10\%$, while Centaur can reach up to $89.90\%$ accuracy. More reduction in uplink data transmission of UCDs can be achieved by a more advanced data selection strategy, lowering the communication burden on constrained devices. This also leads to a notable decrease in overall communication cost compared to AP training, as less data is exchanged between UCDs and APs while improving model accuracy. We remark that model training consumes much more energy and causes much more latency than communication.
Figure 9. The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on MobileNetv3 with CIFAR10 (annotated texts are examples of the classes present on a UCD)
Figure 10. The impacts of imbalanced data partitioning on test accuracy of AP training, UCD training, and Centaur, on ConvNet and UCIHAR dataset dataset.
Correlation among Cost Metrics. To show the connection between cost metrics, we plot MAC, Latency, and Energy in Figure 8. We observe that both Latency and Energy have linear relationships with MAC. Besides, Latency is also linearly correlated with Energy, in such a way that these two entirely overlap when we appropriately scale the y-axes. We remark that Bandwidth also has a similar correlation. Such linear relationships are because of the assumption we made when computing Latency and Energy; however, the actual cost and the relationship among these metrics might not deviate much in practice. Based on such correlations, we can add Latency (or Energy) of model training as well as the communication Bandwidth to get the overall workload. The results shown in Figure 1 indicate that the workload still follows a similar pattern as the computation in model training, because the training workload significantly overweight the communication workload.
6.4. Data and Participation Heterogeneity
Centaur is robust to data heterogeneity. To create imbalanced non-IID data partitions among clients, we use LDA as defined in § 6.1. Next, we set different values for LDA-Alpha to manipulate the levels of non-IID data partitions. Figure 10 and Figure 10 show the results when LDA-Alpha are changing from $0.001$, to $0.1$, to $1$, and to $1000$. The smaller LDA-Alpha is, the less balanced the dataset is. We also annotate some examples of class distribution in Figure 10, demonstrating that LDA-Alpha=0.001=0.001 generates almost $1$ class per client, while LDA-Alpha=1000=1000 generates almost uniform classes per client. Results show that in all cases Centaur can reach higher test accuracy than both AP and UCD training. We remark that due to the retraining of the encoder, both Centaur and AP training can cause more fluctuations than UCD training; this is obvious when data distribution tends to be IID (e.g., when LDA-Alpha=1000=1000 in Figure 10).
Centaur scales well with higher total and participating clients. Here, we scale the FL training problem with the number of total clients from 10 to 1000. $10\%$, $20\%$, and $50\%$ of them are selected respectively as participating clients in each round. Figure 11 shows Centaur achieves higher test accuracy compared to both UCD training and AP training. Specifically, with a higher participation rate, all training methods tend to have better performance. However, with a much larger number of total clients (e.g., 1000), the test accuracy is reduced, because the same size of data is partitioned into more portions. Also, with such more fragmented data partitioning, Centaur has much accuracy gain when compared to UCD training. This experiment focuses on test accuracy and does not account for the orchestration costs of running Centaur in dynamic mobile environments. However, such scalability challenges are not unique to Centaur but are inherent to any FL framework. We also note that practical deployment at larger scales necessitates additional server-side coordination. Centaur is designed to be modular, enabling integration into existing FL stacks. Nevertheless, evaluating the end-to-end overhead of managing varying participation rates and asynchronous aggregation remains open.
6.5. Performance of Centaur under Mobility
Figure 11. The impact of participating vs. total clients on test accuracy for MobileNetv3 with CIFAR10.
Figure 12. The performance of access point (AP) training, usage-constrained device (UCD) training, and partition-based federated learning (Centaur) under different connection probabilities $\lambda$, reported on MobileNetv3 and CIFAR10.
We study when the connection probability $\lambda$ changes based on the mobility model defined in §6.1. For a more readable visualization, we bucket $\lambda$ in three different ranges of $\lambda\in$ [0.1,0.4], $\lambda\in$ [0.4,0.7], and $\lambda\in$ [0.7,1.0] and report the corresponding accuracy when the UCDs are mobile. Figure 12 shows Centaur always has higher efficiency than standard UCD training while incurring a higher latency than direct AP training. Specifically, Centaur gains $\%, 5.645.64\%$, and $4.11\%$ higher accuracy with $51.73\%$, $60.48\%$, and $68.35\%$ lower cost both in terms of energy and latency across the different ranges of $\lambda$. Interestingly, we also observe a rise in accuracy with lower connectivity probability. The main reason is the design of our experimental setup: we realistically assume that with limited connectivity, UCDs devices can collect more data and perform data selection locally. Using this selected data, each device then trains the classifier for more epochs, which further improves its accuracy.
7. Discussion
While Centaur shows strong empirical performance and feasibility on constrained devices, some of our assumptions need further study. We outline key limitations, and opportunities for future work, below.
(1) Centaur assumes each UCD has a trusted, resourceful personal AP (e.g., a smartphone or router) that is always available. This simplifies privacy and connectivity but may not hold in transient or multi-user settings.
(2) We used RaspberryPi as a reproducible proxy for constrained devices, though it is more capable than many real UCDs. Results may thus overestimate feasibility; testing on truly constrained hardware is needed to validate our claims.
(3) Freezing the encoder and training only the classifier reduces on-device compute and memory but limits adaptability: if local data diverge from the pretrained one, classifier-only updates may be insufficient. Studying trade-offs in encoder vs. classifier update frequency remains open.
(4) Our on-device selection uses loss and gradient norm for importance sampling, but this remains heuristic without formal convergence or variance guarantees. It may also be sensitive to outliers, mislabeled, or out-of-distribution data, potentially biasing updates.
(5) Centaur reduces on-device workload but increases AP computation and network use. We did not quantify orchestration and scaling costs with many asynchronous users. Although fewer samples are sent than baselines, uplink and energy costs for constrained devices and overall communication efficiency need further analysis.
(6) We use FedAvg in our implementation, but Centaur can accommodate other FL aggregation strategies. Layer-wise FL methods that perform model training in a partitioned manner, such as FedMA wang2020federated, can also be integrated. Also, we evaluated Centaur on supervised tasks, yet a complementary approach can be suggested for unsupervised tasks.
8. Conclusions
We propose an efficient training method that leverages data selection to improve learning and reduce computational costs in multidevice federated learning. Our approach allows partition-based training and aggregation on resource-constrained devices, supported by resourceful companion devices. Our evaluations over various scenarios confirm that our proposed solution improves model accuracy while reducing system overhead.